Assuming the exponential failure distribution, calculate the probability of a system surviving an operation time equal to twice the duration of the MTBF.
Answer #1

MTBF =1/ λ

P(X > 2 * MTBF)

= P(X > 2/ λ )

=e^(- λ * 2/ λ )

= e^(-2)

= 0.135335